OpenStudy (honeymonster):

please help me solve this linear equation..fine the value of x..equation posted in replies

OpenStudy (honeymonster):

\[\frac{ 4x }{ 5 }=-2+\frac{ 2x }{ 3 }\]

OpenStudy (ajprincess):

1.Add 2 both sides. 2.Multiply by 15 both sides. can u tell me vat u get?

OpenStudy (honeymonster):

so far i have \[\frac{ 4x }{ 5 }+2 = \frac{ 2x }{ 3 }\]

OpenStudy (ajprincess):

\(\large15(\frac {4x}{5}+2)=15(\frac{2x}{3})\)

OpenStudy (ajprincess):

vat do u get nw?

OpenStudy (honeymonster):

err..i think i've calculated it wrong because i got 42 and 10 :/

OpenStudy (ajprincess):

 Nw vat do u get when u subtract 30 both sides?